find the Marshallian demand funtions for commodity x and y


1
Expert's answer
2021-05-05T14:06:41-0400

It deals with utility approach in which consumer maximizes his utility subject to budget constraint. For this case, there is two goods x and y.

If the utility function U(x,y)=xy

Budget constraint

"P_x \times X+P_y\times y =m"

The graph for the Marshallian demand functions is given below -